Fresno, Inc. has monthly sales of 280 units with a sales price of $38 per unit and a variable cost per unit of $16. Currently, the firm has a cash only policy. Jim, the firm’s financial manager, estimates the firm can sell an additional 25 units per month if the firm would switch to a net 30 credit policy. The monthly interest rate is .3 percent. What is the net present value of the switch if you ignore taxes?

$200,333

$172,293

$194,373

$164,033
Answer #1

Incremental cash inflow = (Sales – variable cost) x No. of additional sales units

                                   = ($ 38 - $ 16) x 25 = $ 22 x 25 = $ 550

Present value of incremental cash inflow = $ 550/0.3 % = $ 550/0.003 = $ 183,333.3333

Cost of switching = ($ 38 x 280) + ($ 16 x 25) = $ 10,640 + $ 400 = $ 11,040

NPV of switching = PV of incremental cash flow – cost of switching

                         = $ 183,333.3333 - $ 11,040 = $ 172,293.3333 or $ 172,293.33

Hence option “ $ 172,293” is correct answer.